# Can Language Models Replace Programmers? RepoCod Says ‘Not Yet’

### v2

@@ Abstract
Large language models (LLMs) have achieved high accuracy, i.e., more than 90 pass@1, in solving Python coding problems in HumanEval and MBPP. Thus, a natural question is, whether LLMs achieve comparable code completion performance compared to human developers? Unfortunately, one cannot answer this question using existing manual crafted or simple (e.g., single-line) code generation benchmarks, since such tasks fail to represent real-world software development tasks. In addition, existing benchmarks often use poor code correctness metrics, providing misleading conclusions.
To address these challenges, we create RepoCod, a code generation benchmark with 980 problems collected from 11 popular real-world projects, with more than 58% of them requiring file-level or repository-level context information. In addition, RepoCod has the longest average canonical solution length (331.6 tokens) and the highest average cyclomatic complexity (9.00) compared to existing benchmarks. Each task in RepoCod includes 313.5 developer-written test cases on average for better correctness evaluation. In our evaluations of ten LLMs, none of the models achieve more than 30 pass@1 on RepoCod, indicating the necessity of building stronger LLMs that can help developers in real-world software development. RepoCod is available at https://github.com/lt-asset/REPOCOD
@@ 1 Introduction
LLMs are ineffective at generating functions with high content complexity (up to 20 pass@1 for tasks requiring repository-level context).
@@ 5.1 SOTA LLMs’ Pass@1 on RepoCod
Table 4 shows ten LLMs’ performance on RepoCod, under three retrieval settings. On all retrieval methods, commercial LLMs have better performance. Specifically, GPT-4o has the best result, reaching up to 27.35 pass@1. On the other hand, none of the open-sourced LLMs has over 20 pass@1 in any retrieval settings.
This result shows that SOTA LLMs still struggle with repository-level code generation. Compared to their pass@1 on HumanEval (about 90 Guo et al. (2024)) and MBPP, SOTA LLMs are still far away from writing real-world programs requiring repository-level information. We provide a detailed discussion of the results below.
@@ 5.2 Performance under Different Complexity
Canonical Solution Length. Figure 6(b) presents the performance of LLMs across different length complexity. Similar to the findings from the previous section, the pass@1 of the LLMs gradually decreases as the complexity of the functions increases. In the challenging scenarios (token length > 232), even the top-performing model, GPT-4o, achieves less than 10 pass@1. This low pass@1 performance indicates that while SOTA LLMs are effective at solving simpler problems, they struggle significantly with problems that have medium to long canonical solution lengths in RepoCod. Combining the fact that RepoCod has the longest canonical solution length (Table 2) and the most problems with repository-level context (Table 3), RepoCod sets a new standard for LLMs in code generation.
@@ 7 Conclusion
We present RepoCod, a real-world, complex dataset designed for code generation tasks with repository-level context. RepoCod comprises 980 instances from 11 popular Python repositories, including 257 requiring repository-level context and 316 requiring file-level context, highlighting the complexity and comprehensiveness of the benchmark. Additionally, we introduce an innovative automatic extraction method that can scale RepoCod to any Python project, offering a scalable solution for repository-level code generation benchmarks. Our evaluation of SOTA LLMs on RepoCod reveals a maximum pass@1 of 27.35, with even lower scores for functions requiring repository-level context, underscoring that current models fall short of generating realistic repository-level code. This work underscores the need for further research in repository-level code completion.
@@ A.2 Evaluation Result By Repository
First, repositories differ significantly in difficulty. Models perform well on flask, with Pass@1 scores often above 40, indicating an alignment with LLM capabilities. Second, GPT4-o is the strongest LLM on RepoCod, consistently outperforming or matching other models across all repositories. Finally, LLMs exhibit specialization across repositories. For instance, while Claude 3.5 Sonnet and DeepSeekCoder-33B perform similarly overall, Claude excels on pylint, whereas DeepSeekCoder performs better on plotly.py.

### v3

@@ Abstract
Large language models (LLMs) have achieved high accuracy, i.e., more than 90% pass@1, in solving Python coding problems in HumanEval and MBPP. Thus, a natural question is, whether LLMs achieve comparable code completion performance compared to human developers? Unfortunately, one cannot answer this question using existing manual crafted or simple (e.g., single-line) code generation benchmarks, since such tasks fail to represent real-world software development tasks. In addition, existing benchmarks often use poor code correctness metrics, providing misleading conclusions.
To address these challenges, we create RepoCod, a code generation benchmark with 980 problems collected from 11 popular real-world projects, with more than 58% of them requiring file-level or repository-level context information. In addition, RepoCod has the longest average canonical solution length (331.6 tokens) and the highest average cyclomatic complexity (9.00) compared to existing benchmarks. Each task in RepoCod includes 313.5 developer-written test cases on average for better correctness evaluation. In our evaluations of ten LLMs, none of the models achieve more than 30% pass@1 on RepoCod, indicating the necessity of building stronger LLMs that can help developers in real-world software development. RepoCod is available at https://github.com/lt-asset/REPOCOD
@@ 1 Introduction
LLMs are ineffective at generating functions with high content complexity (up to 25% pass@1 for tasks requiring repository-level context).
@@ 5.1 SOTA LLMs’ Pass@1 on RepoCod
Table 4 shows ten LLMs’ performance on RepoCod, under three retrieval settings. On all retrieval methods, commercial LLMs have better performance. Specifically, GPT-4o has the best result, reaching up to 27.35% pass@1. On the other hand, none of the open-sourced LLMs has over 20% pass@1 in any retrieval settings.
This result shows that SOTA LLMs still struggle with repository-level code generation. Compared to their pass@1 on HumanEval (about 90% Guo et al. (2024)) and MBPP, SOTA LLMs are still far away from writing real-world programs requiring repository-level information. We provide a detailed discussion of the results below.
@@ 5.2 Performance under Different Complexity
Canonical Solution Length. Figure 6(b) presents the performance of LLMs across different length complexity. Similar to the findings from the previous section, the pass@1 of the LLMs gradually decreases as the complexity of the functions increases. In the challenging scenarios (token length > 232), even the top-performing model, GPT-4o, achieves less than 10% pass@1. This low pass@1 performance indicates that while SOTA LLMs are effective at solving simpler problems, they struggle significantly with problems that have medium to long canonical solution lengths in RepoCod. Combining the fact that RepoCod has the longest canonical solution length (Table 2) and the most problems with repository-level context (Table 3), RepoCod sets a new standard for LLMs in code generation.
@@ 7 Conclusion
We present RepoCod, a real-world, complex dataset designed for code generation tasks with repository-level context. RepoCod comprises 980 instances from 11 popular Python repositories, including 257 requiring repository-level context and 316 requiring file-level context, highlighting the complexity and comprehensiveness of the benchmark. Additionally, we introduce an innovative automatic extraction method that can scale RepoCod to any Python project, offering a scalable solution for repository-level code generation benchmarks. Our evaluation of SOTA LLMs on RepoCod reveals a maximum pass@1 of 27.35%, with even lower scores for functions requiring repository-level context, underscoring that current models fall short of generating realistic repository-level code. This work underscores the need for further research in repository-level code completion.
@@ A.2 Evaluation Result By Repository
First, repositories differ significantly in difficulty. Models perform well on flask, with pass@1 often above 40% , indicating an alignment with LLM capabilities. Second, GPT4-o is the strongest LLM on RepoCod, consistently outperforming or matching other models across all repositories. Finally, LLMs exhibit specialization across repositories. For instance, while Claude 3.5 Sonnet and DeepSeekCoder-33B perform similarly overall, Claude excels on pylint, whereas DeepSeekCoder performs better on plotly.py.